OpenStudy (anonymous):

Let f(x)=5x+12. Find f1(x).

OpenStudy (tkhunny):

Do you mean \(f^{-1}(x)\)? Usually, "Swap and Solve". Given y = 5x+12 Swap x and y ==> x = 5y+12 Solve for y

OpenStudy (anonymous):

idk what it means that is just what the question says

OpenStudy (tkhunny):

Nonresponsive. Look VERY CAREFULLY. It's not acceptable for you not to recognize the notation. You have to understand what you are seeing or you have no chance to succeed. Is there the "f1" as you have written, or this \(f^{-1}\) as I have guessed?

OpenStudy (anonymous):

it says find \[f ^{1}(x)\]

OpenStudy (tkhunny):

Awesome. Since that doesn't necessarily mean anything, without additional information, we cannot solve this problem. If there is a negative sign in there, and you're just not sharing, THEN we can fix the question and solve the problem. Otherwise, we're stuck in a missing definition.

OpenStudy (anonymous):

i see the negative sign now it is just hard to see because it looks like the f is capitalized because they have it so close together

OpenStudy (anonymous):

y= 5x +12, solve for x... (y-12)/5=x...the domain of the nverse is the range of function so switch the x and y around y=(x-12)/5 = f-1(x)

OpenStudy (anonymous):

i dont get what you are saying

OpenStudy (tkhunny):

Just swap and solve. Think about other stuff when you are done with that algebra.